Sarah's gas meter readings were 5,003 CCF last month and 5,014 CCF this month. Her natural gas supplier charges $1.72 per CCF. What is her total natural gas bill for the month?



1
Expert's answer
2022-03-23T18:28:24-0400

The CFF of natural gas used = The current reading subtract Previous reading

Therefore

The CCF of natural gas they used "= 5014-5003=11CCF"

"1CCF\to \$1.72"

"\therefore 11CCF\to 11\times 1.72=\$18.92"